Show that the extremum value of the function  f(x) = 1 / (sinx+4)  - 1 / (cos x -4), where x belongs to R is   2√2 / (4 √2 + 1).

Grade:12

1 Answers

Ramesh V
14 years ago

f '(x) = -cos x/(4+sin x)2 -sin x/(-4 +cos x)2

sin x.(4+sin x)2 +cos x.(-4 +cos x)2 = 0

(cos x + sin x)(8sin x -8 cos x + cos x.sin x +17) = 0

as (8sin x -8 cos x + cos x.sin x +17) cant be zero since  -1<sinx<1 and -1<cosx<1

so, (cos x + sin x) = 0

which mean cos x = -sin x

x = 2nPi + 3Pi/4

one possible value is 135o

on substitution, we get

f(x) = 1/(4+sin x) -1/(-4 +cos x)

f(x) = 2/(4+sin x)

put sin x = 2-1/2

which gives extreme value of f(x) as 2√2 / (4 √2 + 1)

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and
we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch